how do I work this problem x-8 over 5 is equal to 2 over 4
Answer by rfer(16322)

(X-8)/5=2/4
WORK IT JUST LIKE YOU WOULD WORK A RATIO
4x-32=10
now, just solve
4x=10+32
4x=42
x=42/4
x=10.5